Broussard Skateboard's sales are expected to increase by 15% from $8.8 million in 2016 to $10.12 million in 2017. Its assets totaled $6 million at the end of 2016. Broussard is already at full capacity, so its assets must grow at the same rate as projected sales. At the end of 2016, current liabilities were $1.4 million, consisting of $450,000 of accounts payable, $500,000 of notes payable, and $450,000 of accruals. The after-tax profit margin is forecasted to be 5%, and the forecasted payout ratio is 65%. What would be the additional funds needed? Do not round intermediate calculations. Round your answer to the nearest dollar.
Answer #1

AFN = Increase in total assets - Increase in liabilities - Addition to retained earnings

Increase in total assets = Assets * Growth rate = 6000000 * 15% = 900000

Spontaneous liabilities = Accounts payable + accruals = 450000 + 450000 = 900000

Increase in spontaneous liabilities = 900000 * growth rate = 900000 * 15% = 135000

Increase in retained earnings = Profit margin * retention ratio * forecasted sales

= 5% * (100% -65%) * 10120000

= 177100

AFN = 900000 - 135000 - 177100

= 587900